OpenStudy (anonymous):

Let A and B be two events in a sample space S such that P(A) = 0.6, P(B) = 0.5, and P(A B) = 0.15. Find the probabilities P(A|B^c) -->(B compliment)

OpenStudy (anonymous):

Can someone give me the generic formula for this too? It always confuses me.

OpenStudy (anonymous):

you can most easily do it with a venn diagram rather than a formula, but \[P(A|B^c)=\frac{P(A\cap B^c)}{P(B^c)}\]

OpenStudy (anonymous):

So would the answer just ne p(A)?

OpenStudy (anonymous):

since \[P(B)=.5\] you know \[P(B^c)=1-.5=.5\]

OpenStudy (anonymous):

(.6 * .5)/.5?

OpenStudy (anonymous):

no no hold on

OpenStudy (anonymous):

you cannot multiply the probabilities together to get the intersection unless they are independent \[P(A\cap B^c)=.45\]

OpenStudy (anonymous):

let me draw a picture to make it clear

OpenStudy (anonymous):

since the probability of the intersection of A and B is .15, and since the probability of A is .6 you get the probability of \(A\cap B^c =.45\)

OpenStudy (anonymous):

so your "final answer" is \[\frac{.45}{.5}=\frac{45}{50}=.9\]

OpenStudy (anonymous):

there was a mistake in my picture, the number outside should have been 5

OpenStudy (anonymous):